# An Engorgio Prompt Makes Large Language Model Babble on

### v1

@@ Abstract
Auto-regressive large language models (LLMs) have yielded impressive performance in many real-world tasks. However, the new paradigm of these LLMs also exposes novel threats. In this paper, we explore their vulnerability to inference cost attacks, where a malicious user crafts Engorgio prompts to intentionally increase the computation cost and latency of the inference process. We design Engorgio, a novel methodology, to efficiently generate adversarial Engorgio prompts to affect the target LLM’s service availability. Engorgio has the following two technical contributions. (1) We employ a parameterized distribution to track LLMs’ prediction trajectory. (2) Targeting the auto-regressive nature of LLMs’ inference process, we propose novel loss functions to stably suppress the appearance of the <EOS> token, whose occurrence will interrupt the LLM’s generation process. We conduct extensive experiments on 13 open-sourced LLMs with parameters ranging from 125M to 30B. The results show that Engorgio prompts can successfully induce LLMs to generate abnormally long outputs (i.e., roughly 2-13$\times$ longer to reach 90%+ of the output length limit) in a white-box scenario and our real-world experiment demonstrates Engergio’s threat to LLM service with limited computing resources. The code is accessible at https://github.com/jianshuod/Engorgio-prompt.
@@ 1 Introduction
Large language models (LLMs) (Touvron et al., 2023a; Ouyang et al., 2022; Carlini et al., 2021) have demonstrated remarkable performance in various real-world applications, e.g., online chatting (Shen et al., 2023), customer service (Gimpel et al., 2023), and finance (Wu et al., 2023). Given the increasing popularity and adoption of LLMs, reducing their inference cost becomes critical. Firstly, from the cost aspect, a modern LLM normally contains billions of parameters, and each inference generation may consume considerable resources and time. Many AI service providers are paying more bills to support their LLM inference services than training (Patel et al., 2024; Li et al., 2024; Patterson et al., 2022). Secondly, from the service availability aspect, there is fierce competition across different LLM service providers, making service reliability and fast response time important factors in attracting customers. Meanwhile, these two considerations motivate malicious entities to attack the LLMs, increasing their operational cost and generation latency.
@@
Figure 1: Distributions of the total lengths (input plus output) of normal samples from ShareGPT (sharegpt, 1) and Engorgio prompts.
In this paper, we introduce Engorgio [^1] Engorgio is a spell in the Harry Potter universe, which causes objects (or creatures) to increase in size., a simple yet effective method to generate threatening Engorgio prompts against state-of-the-art LLMs. Our focus is on text completion, where LLMs predict the next token based on the initial prompt and previously generated tokens until an end-of-sequence (<EOS>) token is predicted or a maximum length is reached. Technically, Engorgio effectively addresses the above challenges via: (1) Inspired by the special role of <EOS> token in determining whether the model halts its response, we adopt an untargeted objective called <EOS> escape loss, which reduces the <EOS> token’s occurrence probability. We also combine a self-mentor loss to stably induce longer responses. (2) We employ a re-parameterization design to effectively utilize the gradients, by modeling the potential distribution of the entire context that can fulfill both objectives. Figure 1 shows the effects of our attack: normal prompts (e.g., the renowned ShareGPT dataset (sharegpt, 1)) typically tempt the LLMs to produce short sequences; in contrast, the crafted Engorgio prompts can make the model response extraordinarily long.
@@ 2.1 Large Language Models (LLMs)
where $P(x_{i}|x_{1},\cdots,x_{i-1})$ denotes the probability of predicting $x_{i}$ as the next token given the sequence $x_{1}\cdots x_{i-1}$. For a Transformer-based model $f_{\Theta}:\mathcal{X}\to\mathcal{Y}$, it accepts a sequence of tokens with any admitted length $S$ and produces an output vector $r_{S}=f_{\Theta}(X_{1:S})\in\mathbb{R}^{V}$ to predict the next token, where $V$ is the model’s vocabulary size. Most prevalent LLMs like LLaMA (Touvron et al., 2023a) and GPT-4 (OpenAI, 2023) are based on the Transformer decoder architecture (Vaswani et al., 2017). The architecture is designed to perform inference in an auto-regressive manner (Graves, 2013), i.e., LLMs generate one token at a time and use the previously generated tokens to instruct the prediction of the next token. We detail the LLM generation process and models involved in this work in Appendix A.1.
@@ 3.1 Attack Insight and Overview
    Figure 2: The pipeline of Engorgio. The whole pipeline is divided into two stages. During the generation stage, we employ a gradient-based method to update the proxy distribution for the Engorgio prompt, where the gradient information is obtained from a local proxy model. For the testing stage, we leverage the optimized proxy distribution to decide the final Engorgio prompt, which is expected to escape the occurrence of the <EOS> token.
@@ 4.1 Experimental Setup
LLMs. We include multiple base models, OPT-125M (Zhang et al., 2022), OPT-1.3B (Zhang et al., 2022), GPT2-large (Radford et al., 2019), LLaMA-7B (Touvron et al., 2023a), LLaMA-2-7B (Touvron et al., 2023b), and LLaMA-30B (Touvron et al., 2023a). SFT models are further fine-tuned with additional dataset, for which we consider seven well-known SFT models including Alpaca (7B) (Taori et al., 2023), Vicuna (7B) (Chiang et al., 2023), StableLM (7B) (stablelm, 2), Koala (7B) (koala, 3), Oraca (7B) (Mukherjee et al., 2023), Samantha (7B) (samantha, 4), and ChatGLM (6B) (Du et al., 2022), as our targets. Considering the crucial importance of prompts, we also consider the three cases of deploying base models with prompts according to the attacker’s knowledge about the deployed prompt. Constrained by space limit, the corresponding results are presented in Appendix B.3.
@@
Baselines. We consider four types of inputs as baselines for comparisons. (1) Normal inputs: we collect 50 samples from the training dataset for Standford-alpaca (Taori et al., 2023), which are generated by OpenAI’s text-davinci-003, and 50 samples from ShareGPT (sharegpt, 1), a website where people can share their ChatGPT conversations. We use the mixup to roughly represent the normal response length of LLMs. (2) Special inputs: we use prompts with the semantics of demanding a longer output ($i.e.$, prompts starting with “output longer”). (3) LLMEffiChecker: we adopt the three attacks (character, word, and structure attack) proposed in Feng et al. (2024) and report the averaged results across the attack variants. (4) Sponge examples: we generate such inputs using the method from Shumailov et al. (2021) by only setting the same input length as our method.
@@
Metrics. Due to the intractable serving mechanisms for LLM, we report results on the level of model behaviors. To mitigate potential sampling bias caused by the inherent variability in LLM inference, we measure the average token number of the generated outputs (Avg-len). We query the target LLM multiple times using the sampling generation and compute the average length across these responses. This renders Avg-len a robust estimate of the Engorgio prompt’s efficacy. Second, we calculate the ratio of the LLM outputs that reach the maximum length (Avg-rate) to evaluate the stability. Notably, inference costs increase super-linearly with longer responses, making Avg-len a lower bound on the prompt’s impact on inference cost. Further analysis of the Engorgio prompt’s impact is detailed in Appendix A.5.
@@
Configurations. We use the Adam optimizer with a learning rate of 0.1 to update the distribution matrix $\theta$. We allow a maximum of 300 optimization steps, the cost of which is acceptable, especially when considering the reusability as explained in Appendix A.6. The Gumbel-Softmax temperature factor $\tau$ is set to 1, and the default Engorgio prompt length is $t=32$. The input length of normal inputs, special inputs, LLMEffiChecker, and sponge examples is roughly the same as Engorgio to ensure fairness. According to our extensive experiments, the loss coefficient $\lambda$ is empirically set to 1. The optimization starts with a random prompt and we use it to initialize the proxy distribution. Constrained by the computing resources, we set 1,024 as the pre-set maximum length. We also conduct experiments with full context size on two representative base models (2,048 for LLaMA-30B and LLaMA-7B) and one SFT model (4,096 for Samantha) to demonstrate the extensibility to longer context size. Please refer to Appendix B.8 for examples of prompts and responses for normal inputs, sponge examples, and Engorgio prompts.
@@ 4.3 Ablation Study
Impact of loss design. We explore the efficacy of the two loss terms. Initially, we assess the performance when optimizing only with the <EOS> escape loss (noted as “ESC” in Table 3). A comparison with normal input and special input from Table 1 reveals that even utilizing only the <EOS> escape loss consistently results in longer outputs. We also observe that combining the self-mentor loss, identified as “ESC+Self-mentor” in Table 3, further increases the Avg-len with almost no extra cost. More experiments in Appendix B.5 show that the performance of Engorgio is not strictly dependent on the choice of $\lambda$.
@@ 4.6 Why is Our Method Effective?
Q1. How does the distribution matrix instruct the token selection in the optimization process?
In the optimization process, we seek to update the distribution matrix $\theta$ rather than selecting individual tokens, meaning that the most suitable input is the aggregated embedding $\tilde{e}(\theta_{i})$. In Figure 5, we find that even in the middle of the optimization process, the token sequence greedily sampled according to the distribution matrix performs only slightly worse than the aggregated embeddings when our goal is to decrease the total loss. This means that our distribution update design is effective in searching for suitable token sequences. Moreover, we find that the <EOS> escape loss of SFT models like Koala is much harder to decrease than base models like OPT-125M. This partially supports that base models are easier to induce than SFT models.
@@
Q2. Does <EOS> escape loss stop <EOS> from appearing?
To verify that <EOS> escape loss reduces the probability of <EOS> appearance, we calculate the highest probability of the <EOS> token at all the $S$ positions. We formulate this as $\mu=\max(\{{\text{Softmax}(f_{\Theta}(E(\theta)_{1:i}))_{\kappa}}\}_{i=1}^{S})$ which signals the highest probability of the interruption of the generation process. We report the relative level of $\mu$ compared to the average probability $1/V$ coupled with the change of <EOS> escape loss in Figure 5. We find that the decrease of <EOS> escape loss can lead the maximum probability of the occurrence of <EOS> token to a low level (close to 0). This substantiates the effectiveness of <EOS> escape loss in stopping <EOS> token from appearing.
@@ A.1 Involved models
Base model GPT-2 (Radford et al., 2019) 2019 117M, 345M, 762M, 1.5B
OPT (Zhang et al., 2022) 2022 125M, 350M, 1.3B, 2.7B, 6.7B, 13B, 66B, 175B
LLaMA (Touvron et al., 2023a) 2023 7B, 13B, 30B, 65B
LLaMA-2 (Touvron et al., 2023b) 2023 7B, 13B, 30B, 65B
  SFT model Alpaca (Taori et al., 2023) LLaMA 2023 7B
Vicuna (Chiang et al., 2023) LLaMA 2023 7B, 13B
Koala (koala, 3) LLaMA 2023 7B, 13B, 30B, 65B
StableLM (stablelm, 2) StableLM-Base 2023 3B, 7B
Orca (Mukherjee et al., 2023) LLaMA-2 2023 7B
Samantha (samantha, 4) LLaMA-2 2023 7B
ChatGLM (Du et al., 2022) ChatGLM-Base 2023 6B, 130B
@@ A.5 Further severity analysis of Engorgio
Figure 7 shows that the output token lengths and the inference costs (FLOPs) of LLaMA-7B (Touvron et al., 2023a) and ChatGLM (Du et al., 2022) are approximately linearly correlated (Kaplan et al., 2020; Hoffmann et al., 2022). It is worth noting that the attention-related operations only account for a small part of the overall operations of the model when $N$ is substantially smaller in magnitude relative to the hidden dimension. (Kaplan et al., 2020; Hoffmann et al., 2022) That’s why we observed an approximately linear plot in Figure 7. The $O(N^{2})$ complexity means an incalculable number of FLOPs when a larger output length is induced, implying the more severe threats of Engorgio to the inference process of the decoder-based models with larger pre-set maximal lengths.
@@ B.1 Transferability for black-box setting
In the limited-knowledge scenario, we evaluate the transferability of Engorgio prompts. We employ a proxy model to craft Engorgio prompts and gauge their impact on target models. Our investigation brings the results as detailed in Table 7. To reason about the results, we have also explored the potential rationales for transferability in Section 7.
Table 7: Results of transferability to different models. Both normal inputs and Engorgio point to the target “To” model.

### v2

@@ Abstract
Auto-regressive large language models (LLMs) have yielded impressive performance in many real-world tasks. However, the new paradigm of these LLMs also exposes novel threats. In this paper, we explore their vulnerability to inference cost attacks, where a malicious user crafts Engorgio prompts to intentionally increase the computation cost and latency of the inference process. We design Engorgio, a novel methodology, to efficiently generate adversarial Engorgio prompts to affect the target LLM’s service availability. Engorgio has the following two technical contributions. (1) We employ a parameterized distribution to track LLMs’ prediction trajectory. (2) Targeting the auto-regressive nature of LLMs’ inference process, we propose novel loss functions to stably suppress the appearance of the <EOS> token, whose occurrence will interrupt the LLM’s generation process. We conduct extensive experiments on 13 open-sourced LLMs with parameters ranging from 125M to 30B. The results show that Engorgio prompts can successfully induce LLMs to generate abnormally long outputs (i.e., roughly 2-13$\times$ longer to reach 90%+ of the output length limit) in a white-box scenario and our real-world experiment demonstrates Engergio’s threat to LLM service with limited computing resources. The code is released at: https://github.com/jianshuod/Engorgio-prompt.
@@ 1 Introduction
Large language models (LLMs) (Touvron et al., 2023; Ouyang et al., 2022; Carlini et al., 2021) have demonstrated remarkable performance in various real-world applications, e.g., online chatting (Shen et al., 2023), customer service (Gimpel et al., 2023), and finance (Wu et al., 2023). Given the increasing popularity and adoption of LLMs, reducing their inference cost becomes critical. Firstly, from the cost aspect, a modern LLM normally contains billions of parameters, and each inference generation may consume considerable resources and time. Many AI service providers are paying more bills to support their LLM inference services than training (Patel et al., 2024; Li et al., 2024; Patterson et al., 2022). Secondly, from the service availability aspect, there is fierce competition across different LLM service providers, making service reliability and fast response time important factors in attracting customers. Meanwhile, these two considerations motivate malicious entities to attack the LLMs, increasing their operational cost and generation latency.
@@
Figure 1: Distributions of the total lengths (input plus output) of normal samples from ShareGPT22footnotemark: 2 and Engorgio prompts.
In this paper, we introduce Engorgio [^1] Engorgio is a spell in the Harry Potter universe, which causes objects (or creatures) to increase in size., a simple yet effective method to generate threatening Engorgio prompts against state-of-the-art LLMs. Our focus is on text completion, where LLMs predict the next token based on the initial prompt and previously generated tokens until an end-of-sequence (<EOS>) token is predicted or a maximum length is reached. Technically, Engorgio effectively addresses the above challenges via: (1) Inspired by the special role of <EOS> token in determining whether the model halts its response, we adopt an untargeted objective called <EOS> escape loss, which reduces the <EOS> token’s occurrence probability. We also combine a self-mentor loss to stably induce longer responses. (2) We employ a re-parameterization design to effectively utilize the gradients, by modeling the potential distribution of the entire context that can fulfill both objectives. Figure 2 shows the effects of our attack: normal prompts (e.g., the renowned ShareGPT dataset [^2] https://sharegpt.com) typically tempt the LLMs to produce short sequences; in contrast, the crafted Engorgio prompts can make the model response extraordinarily long.
@@ 2.1 Large Language Models (LLMs)
where $P(x_{i}|x_{1},\cdots,x_{i-1})$ denotes the probability of predicting $x_{i}$ as the next token given the sequence $x_{1}\cdots x_{i-1}$. For a Transformer-based model $f_{\Theta}:\mathcal{X}\to\mathcal{Y}$, it accepts a sequence of tokens with any admitted length $S$ and produces an output vector $r_{S}=f_{\Theta}(X_{1:S})\in\mathbb{R}^{V}$ to predict the next token, where $V$ is the model’s vocabulary size. Most prevalent LLMs like LLaMA (Touvron et al., 2023) and GPT-4 (OpenAI, 2023) are based on the Transformer decoder architecture (Vaswani et al., 2017). The architecture is designed to perform inference in an auto-regressive manner (Graves, 2013), i.e., LLMs generate one token at a time and use the previously generated tokens to predict next tokens. We detail the LLM generation process and models involved in this work in Appendix A.1.
@@ 3.1 Attack Insight and Overview
    Figure 2: The pipeline of Engorgio. The whole pipeline is divided into two stages. During the generation stage, we employ a gradient-based method to update the proxy distribution for the Engorgio prompt, where the gradient information is obtained from a local proxy model. For the testing stage, we leverage the optimized proxy distribution to decide the final Engorgio prompt.
@@ 4.1 Experimental Setup
LLMs. We include multiple base models, OPT-125M, OPT-1.3B, GPT2-large, LLaMA-7B, LLaMA-2-7B, and LLaMA-30B. SFT models are further fine-tuned with additional datasets, for which we consider seven well-known SFT models including Alpaca (7B), Vicuna (7B), StableLM (7B), Koala (7B), Oraca (7B), Samantha (7B), and ChatGLM (6B), as our targets. More details about the models involved in this work are provided in Appendix A.1. Considering the crucial importance of prompts, we also consider the three cases of deploying base models with prompts according to the attacker’s knowledge about the deployed prompt (cf. Appendix B.3).
@@
Baselines. We consider four types of inputs as baselines for comparisons. (1) Normal inputs: we collect 50 samples from the training dataset for Standford-alpaca [^3] https://github.com/tatsu-lab/stanford_alpaca/, which are generated by OpenAI’s text-davinci-003, and 50 samples from ShareGPT [^4] https://sharegpt.com/, a website where people can share their ChatGPT conversations. We use the mixup to roughly represent the normal response length of LLMs. (2) Special inputs: we use prompts with the semantics of demanding a longer output ($i.e.$, prompts starting with “output longer”). (3) LLMEffiChecker: we adopt the three attacks (character, word, and structure attack) proposed in Feng et al. (2024) and report the averaged results across the attack variants. (4) Sponge examples: we generate such inputs using the method from Shumailov et al. (2021) by only setting the same input length as our method.
@@
Metrics. Due to the intractable serving mechanisms for LLM, we report results on the level of model behaviors. To mitigate potential sampling bias caused by the inherent variability in LLM inference, we measure the average token number of the generated outputs (Avg-len). We query the target LLM multiple times using the sampling generation and compute the average length across these responses. This renders Avg-len a robust estimate of the Engorgio prompt’s efficacy. Second, we calculate the ratio of the LLM outputs that reach the maximum length (Avg-rate) to evaluate the stability. Notably, inference costs increase super-linearly with longer responses, making Avg-len a lower bound on the prompt’s impact on inference cost, which we detail in Appendix A.5.
@@
Configurations. We use the Adam optimizer with a learning rate of 0.1 to update the distribution matrix $\theta$. We allow a maximum of 300 optimization steps, the cost of which is acceptable, especially when considering the reusability as explained in Appendix A.6. The Gumbel-Softmax temperature factor $\tau$ is set to 1, and the default Engorgio prompt length is $t=32$. The input length of normal inputs, special inputs, LLMEffiChecker, and sponge examples is roughly the same as Engorgio to ensure fairness. The loss coefficient $\lambda$ is empirically set to 1. The optimization starts with a random prompt, which we use to initialize the proxy distribution. Constrained by the computing resources, we set 1,024 as the pre-set maximum length. We also conduct experiments with full context size on two representative base models (2,048 for LLaMA-30B and LLaMA-7B) and one SFT model (4,096 for Samantha) to demonstrate the extensibility to longer context size. Please refer to Appendix B.8 for examples of prompts and responses for normal inputs, sponge examples, and Engorgio prompts.
@@ 4.3 Ablation Study
Impact of loss design. Initially, we assess the performance when optimizing only with the <EOS> escape loss (noted as “ESC” in Table 3). A comparison with normal input and special input from Table 1 reveals that even utilizing only the <EOS> escape loss consistently results in longer outputs. We also observe that combining the self-mentor loss, identified as “ESC+Self-mentor” in Table 3, further increases the Avg-len with almost no extra cost. More experiments in Appendix B.5 show that Engorgio is not strongly dependent on the choice of $\lambda$.
@@ 4.6 Why is Our Method Effective?
Q1. How does the distribution matrix instruct the token selection in the optimization process? In the optimization process, we seek to update the distribution matrix $\theta$ rather than selecting individual tokens, meaning that the most suitable input is the aggregated embedding $\tilde{e}(\theta_{i})$. In Figure 5, we find that even in the middle of the optimization process, the token sequence greedily sampled according to the distribution matrix performs only slightly worse than the aggregated embeddings when our goal is to decrease the total loss. This means that our distribution update design is effective in searching for suitable token sequences. Moreover, we find that the <EOS> escape loss of SFT models like Koala is much harder to decrease than base models like OPT-125M. This partially supports that base models are easier to induce than SFT models.
@@
Q2. Does <EOS> escape loss stop <EOS> from appearing? To verify that <EOS> escape loss reduces the probability of <EOS> appearance, we calculate the highest probability of the <EOS> token at all the $S$ positions. We formulate this as $\mu=\max(\{{\text{Softmax}(f_{\Theta}(E(\theta)_{1:i}))_{\kappa}}\}_{i=1}^{S})$ which signals the highest probability of the interruption of the generation process. We report the relative level of $\mu$ compared to the average probability $1/V$ coupled with the change of <EOS> escape loss in Figure 5. We find that the decrease of <EOS> escape loss can lead the maximum probability of the occurrence of <EOS> token to a low level (close to 0). This substantiates the effectiveness of <EOS> escape loss in stopping <EOS> token from appearing.
@@ 6 Conclusion
Ethics Statement
This paper highlights potential adversarial threats to LLM service availability. Instead of conducting real-world attacks, this work serves as a clarion call for service providers to consider not only maximizing service latency but also the risks to inference costs posed by malicious users. Engorgio offers a method for generating threatening prompts, allowing service providers to stress test their online LLM services effectively. All experiments adhere to principles of trustworthiness and harmlessness. Note that real-world attack demos in Section 4.5 target only our own LLM service, without impacting others. Our work utilizes open-source models and datasets, ensuring no privacy violations. Our work also does not involve any human subject. This work does not raise ethical issues in general.
Reproducibility Statement
The details of models, hyper-parameter settings, and experimental settings can be found in Section 4.1 and Appendix B.6. The models involved in this work are all openly accessible. The codes for reproducing our main evaluation results are provided in the anonymous repository. We will release the full codes of our methods upon the acceptance of this paper.
Acknowledgments
We would like to thank Yiming Li and the anonymous reviewers for their helpful comments. This work is supported by the National Science Foundation for Distinguished Young Scholars of China under No. 62425201, and the National Science Foundation China under Grant No. 62132011.
@@ A.1 Involved models
  Base model GPT-2666https://github.com/openai/gpt-2  –  2019  117M, 345M, 762M, 1.5B
OPT [^7] https://github.com/facebookresearch/metaseq/  –  2022 125M, 350M, 1.3B, 2.7B, 6.7B, 13B, 66B, 175B
LLaMA [^8] https://ai.meta.com/blog/large-language-model-llama-meta-ai/  –  2023 7B, 13B, 30B, 65B
LLaMA-2999https://www.llama.com/llama2/  –  2023 7B, 13B, 30B, 65B
  SFT model Alpaca [^10] https://github.com/tatsu-lab/stanford_alpaca  LLaMA  2023 7B
Vicuna [^11] https://lmsys.org/blog/2023-03-30-vicuna/  LLaMA  2023 7B, 13B
Koala [^12] https://bair.berkeley.edu/blog/2023/04/03/koala/  LLaMA  2023 7B, 13B, 30B, 65B
StableLM [^13] https://github.com/Stability-AI/StableLM  StableLM-Base  2023  3B, 7B
Orca [^14] https://huggingface.co/pankajmathur/orca_mini_v3_7b  LLaMA-2  2023 7B
Samantha [^15] https://huggingface.co/cognitivecomputations/Samantha-1.11-7b  LLaMA-2  2023 7B
ChatGLM [^16] https://github.com/THUDM/ChatGLM-6B  ChatGLM-Base  2023 6B, 130B
@@ A.5 Further severity analysis of Engorgio
Figure 7 shows that the output token lengths and the inference costs (FLOPs) of LLaMA-7B (Touvron et al., 2023) and ChatGLM (Du et al., 2022) are approximately linearly correlated (Kaplan et al., 2020; Hoffmann et al., 2022). It is worth noting that the attention-related operations only account for a small part of the overall operations of the model when $N$ is substantially smaller in magnitude relative to the hidden dimension. (Kaplan et al., 2020; Hoffmann et al., 2022) That’s why we observed an approximately linear plot in Figure 7. The $O(N^{2})$ complexity means an incalculable number of FLOPs when a larger output length is induced, implying the more severe threats of Engorgio to the inference process of the decoder-based models with larger pre-set maximal lengths.
@@ B.1 Transferability for black-box setting
In the limited-knowledge scenario, we evaluate the transferability of Engorgio prompts. We employ a proxy model to craft Engorgio prompts and gauge their impact on target models. Our investigation brings the results as detailed in Table 7. To explain the results, we have also explored the potential rationales for transferability in this section.
Table 7: Transferability: both normal inputs and Engorgio point to the target “To” model.